# PhyT2V: LLM-Guided Iterative Self-Refinement for Physics-Grounded Text-to-Video Generation

### v1

@@ Abstract
Text-to-video (T2V) generation has been recently enabled by transformer-based diffusion models, but current T2V models lack capabilities in adhering to the real-world common knowledge and physical rules, due to their limited understanding of physical realism and deficiency in temporal modeling. Existing solutions are either data-driven or require extra model inputs, but cannot be generalizable to out-of-distribution domains. In this paper, we present PhyT2V, a new data-independent T2V technique that expands the current T2V model’s capability of video generation to out-of-distribution domains, by enabling chain-of-thought and step-back reasoning in T2V prompting. Our experiments show that PhyT2V improves existing T2V models’ adherence to real-world physical rules by 2.3x, and achieves 35% improvement compared to T2V prompt enhancers. The source codes are available at: https://github.com/pittisl/PhyT2V.
Figure 1: Left: videos generated by the current text-to-video generation model (CogVideoX-5B [46]) cannot adhere to the real-world physical rules (described in brackets following the user prompt). Right: our method PhyT2V, when applied to the same model, better reflects the real-world physical knowledge.
@@ 1 Introduction
Text-to-video (T2V) generation has recently marked a significant breakthrough of generative AI, with the advent of transformer-based diffusion models such as Sora [3], Pika [13] and CogVideoX [47] that can produce videos conditioned on textual prompts. These models demonstrate astonishing capabilities of generating complex and photorealistic scenes, and could even make it difficult for humans to distinguish between real-world and AI-generated videos, in the aspect of individual video frames’ quality [33; 1].
@@
On the other hand, as shown in Figure 1 - Left, current T2V models still have significant drawbacks in adhering to the real-world common knowledge and physical rules, such as quantity, material, fluid dynamics, gravity, motion, collision and causality, and such limitations fundamentally prevent current T2V models from being used for real-world simulation [7; 27; 15]. Enforcement of real-world knowledge and physical rules in T2V generation, however, is challenging because it requires the models’ understandings of not only individual objects but also how these objects move and interact with each other. Further, unlike generating static images, T2V generation requires frame-to-frame consistency in object appearance, shape, motion, lighting and other dynamics [11]. Current T2V models often lack such temporal modeling, especially over long sequences [16], and the generated videos often contain flickering, inconsistent motion and object deformations across frames [22].
@@
Most of the existing solutions to these challenges are data-driven, by using large multimodal T2V datasets that cover different real-world domains to train the diffusion model [45; 10; 37]. However, these solutions heavily rely on the volume, quality and diversity of datasets [47; 38]. Since real-world common knowledge and physical rules are not explicitly embedded in the T2V generation process, the quality of video generation would largely drop in out-of-distribution domains that are not covered by the training dataset, and the generalizability of T2V models is limited due to the vast diversity of real-world scenario domains. Alternatively, researchers also use the existing 3D engines (e.g, Blender [8], Unity3D [32] and Unreal [12]) or mathematical models of edge and depth maps [23; 24; 22] to inject real-world physical knowledge into the T2V model, but these approaches are limited to fixed physical categories and patterns such as predefined objects and movements [45; 22], similarly lacking generalizability.
@@
To achieve generalizable enforcement of physics-grounded T2V generation, we propose a fundamentally different approach: instead of expanding the training dataset or further complicating the T2V model architecture, we aim to expand the current T2V model’s capability of video generation from in-distribution to out-of-distribution domains, by embedding real-world knowledge and physical rules into the text prompts with sufficient and appropriate contexts. To avoid ambiguous and unexplainable prompt engineering [9; 29; 28], our basic idea is to enable chain-of-thought (CoT) and step-back reasoning in T2V generation prompting, to ensure that T2V models follow correct physical dynamics and inter-frame consistency by applying step-by-step guidance and iterative refinement.
@@
We evaluated PhyT2V by applying it onto multiple SOTA T2V models, by using ChatGPT4 o1-preview [14] for LLM reasoning and Tarsier [35] as the video captioning model. We used two major T2V prompt datasets that cover 7 different real-world domains, and compared PhyT2V with the most competitive baselines of prompt enhancers. Our main findings are as follows.
@@ 2.1 T2V Generation Models
Early T2V techniques generate video frames from text-to-image model outputs with temporal extensions [31], but cannot maintain temporal consistency and coherence over time, often producing visually appealing but temporally disconnected outputs. Diffusion Transformers (DiT) [26] improved such consistency with a transformer backbone capable of capturing more complex temporal dynamics and relationships across frames through attention mechanism and long-range dependency modeling [47; 38]. Based on the DiT architecture, recent T2V models, such as OpenSora [49] and VideoCrafter [4], demonstrated that T2V generation can be further improved by in-context learning when provided with sufficient contextual information [40].
@@
However, as shown in Figure 3, although these T2V models demonstrate strong capabilities in video generation when dealing with prompts aligned with the distributions found in the training data, they encounter significant challenges with out-of-distribution prompts that are not covered by training data [^1] In Figure 3, the in-distribution prompts are picked from the ones listed in [46], and the out-of-distribution prompts are our crafted ones for similar scenarios as the in-distribution prompts.. In such cases, the outputs often contain physical illusions or artifacts, reflecting the model’s limitations in generating realistic and coherent video contents under unfamiliar conditions. Such limitations can be addressed by enlarging the training datasets, improving T2V model architectures or developing new mechanisms for adaptation and error correction [41; 39], but these approaches are all prompt-specific and lack generalizability.
@@
On the other hand, as shown in Figure 4, recent research has demonstrated that the quality of video generation with an out-of-distribution prompt can be improved by refining the prompt with sufficient and appropriate details [47; 11]. These findings motivate our design of PhyT2V that embeds contexts of real-world knowledge and physical rules into T2V prompts, to guide the T2V process for better physical accuracy and temporal alignment. The existing works, however, could still fail when tackling more intricate scenarios such as multi-object interactions, because the T2V model lacks an efficient feedback mechanism to learn how the generated video deviates from the real-world knowledge and physical rules. Researchers suggest to provide such feedback with extra input modalities to T2V models such as sampled video frames, depth map or scribble images [40; 48], but incur significant amounts of extra computing overhead and cannot be generalizable. Instead, in our design of PhyT2V, we aim to fully automate the feedback with only text prompts, and enable iterative feedback for the optimum video quality.
@@ 2.2 Using LLM in T2V Generation
LLMs with strong capabilities in natural language processing (NLP) have become a natural choice for prompt refinement in text-to-image and text-to-video generation, and existing work has utilized LLMs to interpret text prompts and orchestrate the initial layout configurations [19; 20; 21; 50; 42; 44]. However, since current LLMs generally lack inherent understandings of the real-world physical laws, using LLMs with simple instructions usually result in videos that appear visually coherent but lack accurate physical realism, particularly when generating scenes with complex object interactions. Furthermore, these approaches frequently rely on static prompts or simple iterative refinements based on bounding box and segmentation map, which may capture basic visual attributes but fail to adapt to nuanced changes that require continuous physical modeling and adjustment.
@@ 3.1 Prompting in Parallel Subproblems for Local CoT Reasoning
Besides, to ensure proper CoT reasoning, we follow the existing work [36; 18] and provide in-context examples $[E]$ about tasks. To facilitate LLM’s in-context learning [5; 6], $[E]$ is given in the format of QA pairs. That is, instead of fine-tuning a separate LLM checkpoint for each new task, we prompt the LLM with a few input-output exemplars, to demonstrate the task and condition the task’s input-output format to the LLM, to guide the LLM’s reasoning process.
@@ 4 Experiments
  • CogVideoX [47]: It can generate 10-second videos aligned from text prompts, with 16 FPS and 768$\times$1360 resolution. It offers two model variants, with 2B and 5B parameters, respectively.
  • OpenSora 1.2 [49]: As an alternative to OpenAI’s Sora model [3], it contains 1.1B parameters and can produce high-quality videos with 16 seconds, 720p resolution and different aspect ratios.
  • VideoCrafter [4]: With 1.8B parameters, it is capable of generating both images and videos from text prompts, at the resolution of 576$\times$1024, with special emphasis on video dynamics.
@@
Baselines: For fair comparison, we only use the existing T2V prompt enhancers as baselines, and other existing work with extra inputs to T2V models [7; 27; 15; 23; 22] are not applicable. We involve two prompt enhancers: 1) Directly using the existing LLM, particularly ChatGPT4, as the prompt enhancer [24; 43]; 2) Promptist [17], which uses reinforcement learning to automatically refine and enhance prompts in the model-preferred way.
@@ 4.2 Different domains of Physical Rules
4.3 Ablation Study
We conduct an ablation study to evaluate the necessity of both the physical rule reasoning (Step 1) and the mismatch reasoning (Step 2) within our PhyT2V workflow, by removing one of these steps from the refinement process to assess its impact on the quality of video generation.
@@ 4.3 Ablation Study
Overall, our study shows that both reasoning steps are integral to the PhyT2V workflow, contributing to a more robust and semantically-aligned generation of refined prompts in Step 3. Detailed ablation studies are in Appendix C.
@@ 5 Conclusion
In this paper, we present PhyT2V, a new T2V technique that expands the existing T2V model’s capability to out-of-distribution domains via LLM reasoning. Experiment results show that PhyT2V can improve the generated video’s adherence to real-world physical rules by up to 2.3x.
@@ Appendix A Details of Evaluation Setup
In our evaluations, we use 4 T2V generation models, including CogVideoX-5B [46], CogVideoX-2B [46], OpenSora [49] and VideoCrafter [4]. We choose to use these models because they are all built with transformer-based diffusion models, which enhance the semantic adherence by using the cross-attention mechanism, and were ranked top on the leaderboards of both the VideoPhy and PhyGenBench benchmarks. We use Tarsier [35] as the video captioning model, because it achieves SOTA performance on multiple video question answering (VQA) datasets, hence ensuring precise and detailed video captioning in our approach.
Since PhyT2V improves the quality of generated videos through text prompt refinement, we use the Promptist [17] and GPT-4o [43] as the prompt enhancers, with the same model hyper-parameter settings as used in PhyT2V, to maintain the consistency between these baselines and PhyT2V.
To fit the maximum token input length of the T2V model, we limit the word length of the refined prompts to 120, by instructing the ChatGPT4 o1-preview model that is used as the LLM for reasoning. To formatting the output and storage in our approach, the ChatGPT4 o1-preview model are instructed to output in JSON format and output results in each step are saved in a CSV file by row. More specifically, in our implementation, we invoke the ChatGPT4 model by using the OpenAI o1 APIs, with our constructed prompts as described in Section 3. In each round of the refinement process, after Step 1 and Step 2 finish, their outputs are embedded to the prompt of Step 3 by replacing the pre-defined place holder, and we then use the generated prompt as the input to LLM again to generated the refined prompt for this round.
@@ Appendix B More Examples of Physics-Grounded Videos Generated by PhyT2V
In this section, we extend Section 4 in the main text of the our paper, and provide more examples of the physics-grounded videos generated by PhyT2V. We will first show more videos generated in different categories of physical rules, and then show more details about the iterative process of prompt refinement.
@@ B.1 Examples of Generated Videos in Different Categories of Physical Rules
In this subsection, we present more examples of generated videos, with the CogVideoX-5B T2V model on both the VideoPhy and PhyGenBench prompt datasets. Fig 11, Fig 12 & 13, and Fig 14 & Figure 15 show the generated videos in categories of fluid-to-fluid, solid-to-fluid, and solid-to-solid for the VideoPhy dataset. Fig 16, 17 and 18 show the generated videos in categories of mechanics, optics and thermal on the PhyGenBench dataset.
B.2 Details of Prompt Refinement in PhyT2V

### v2

@@ Abstract
Text-to-video (T2V) generation has been recently enabled by transformer-based diffusion models, but current T2V models lack capabilities in adhering to the real-world common knowledge and physical rules, due to their limited understanding of physical realism and deficiency in temporal modeling. Existing solutions are either data-driven or require extra model inputs, but cannot be generalizable to out-of-distribution domains. In this paper, we present PhyT2V, a new data-independent T2V technique that expands the current T2V model’s capability of video generation to out-of-distribution domains, by enabling chain-of-thought and step-back reasoning in T2V prompting. Our experiments show that PhyT2V improves existing T2V models’ adherence to real-world physical rules by 2.3x, and achieves 35% improvement compared to T2V prompt enhancers.
@@ 1 Introduction
Text-to-video (T2V) generation has recently marked a significant breakthrough of generative AI, with the advent of transformer-based diffusion models such as Sora [3], Pika [17] and CogVideoX [51] that can produce videos conditioned on textual prompts. These models demonstrate astonishing capabilities of generating complex and photorealistic scenes, and could even make it difficult for humans to distinguish between real-world and AI-generated videos, in the aspect of individual video frames’ quality [37, 1].
@@
On the other hand, as shown in Figure LABEL:fig:coverdemo - Left, current T2V models still have significant drawbacks in adhering to the real-world common knowledge and physical rules, such as quantity, material, fluid dynamics, gravity, motion, collision and causality, and such limitations fundamentally prevent current T2V models from being used for real-world simulation [7, 31, 19]. Enforcement of real-world knowledge and physical rules in T2V generation, however, is challenging because it requires the models’ understandings of not only individual objects but also how these objects move and interact with each other. Further, unlike generating static images, T2V generation requires frame-to-frame consistency in object appearance, shape, motion, lighting and other dynamics [11]. Current T2V models often lack such temporal modeling, especially over long sequences [20], and the generated videos often contain flickering, inconsistent motion and object deformations across frames [26].
@@
Most of the existing solutions to these challenges are data-driven, by using large multimodal T2V datasets that cover different real-world domains to train the diffusion model [49, 10, 41, 12]. However, these solutions heavily rely on the volume, quality and diversity of datasets [51, 42]. Since real-world common knowledge and physical rules are not explicitly embedded in the T2V generation process, the quality of video generation would largely drop in out-of-distribution domains that are not covered by the training dataset, and the generalizability of T2V models is limited due to the vast diversity of real-world scenario domains. Alternatively, researchers also use the existing 3D engines (e.g, Blender [8], Unity3D [36] and Unreal [16]) or mathematical models of edge and depth maps [27, 28, 26] to inject real-world physical knowledge into the T2V model, but these approaches are limited to fixed physical categories and patterns such as predefined objects and movements [49, 26], similarly lacking generalizability.
@@
To achieve generalizable enforcement of physics-grounded T2V generation, we propose a fundamentally different approach: instead of expanding the training dataset or further complicating the T2V model architecture, we aim to expand the current T2V model’s capability of video generation from in-distribution to out-of-distribution domains, by embedding real-world knowledge and physical rules into the text prompts with sufficient and appropriate contexts. To avoid ambiguous and unexplainable prompt engineering [9, 33, 32], our basic idea is to enable chain-of-thought (CoT) and step-back reasoning in T2V generation prompting, to ensure that T2V models follow correct physical dynamics and inter-frame consistency by applying step-by-step guidance and iterative refinement.
@@
For physcial-grounded video generation performance, we further evaluated PhyT2V by applying it onto multiple SOTA T2V models, by using ChatGPT4 o1-preview [18] for LLM reasoning and Tarsier [39] as the video captioning model. We used two major T2V prompt datasets that cover 7 different real-world domains, and compared PhyT2V with the most competitive baselines of prompt enhancers. Our main findings are as follows:
@@ 2.1 T2V Generation Models
Early T2V techniques generate video frames from text-to-image model outputs with temporal extensions [35], but cannot maintain temporal consistency and coherence over time, often producing visually appealing but temporally disconnected outputs. Diffusion Transformers (DiT) [30] improved such consistency with a transformer backbone capable of capturing more complex temporal dynamics and relationships across frames through attention mechanism and long-range dependency modeling [51, 42]. Based on the DiT architecture, recent T2V models, such as OpenSora [53] and VideoCrafter [4], demonstrated that T2V generation can be further improved by in-context learning when provided with sufficient contextual information [44].
@@
However, as shown in Figure 2, although these T2V models demonstrate strong capabilities in video generation when dealing with prompts aligned with the distributions found in the training data, they encounter significant challenges with out-of-distribution prompts that are not covered by training data [^1] In Figure 2, the in-distribution prompts are picked from the ones listed in [50], and the out-of-distribution prompts are our crafted ones for similar scenarios as the in-distribution prompts.. In such cases, the outputs often contain physical illusions or artifacts, reflecting the model’s limitations in generating realistic and coherent video contents under unfamiliar conditions. Such limitations can be addressed by enlarging the training datasets, improving T2V model architectures or developing new mechanisms for adaptation and error correction [45, 43], but these approaches are all prompt-specific and lack generalizability.
@@
On the other hand, as shown in Figure 3, recent research has demonstrated that the quality of video generation with an out-of-distribution prompt can be improved by refining the prompt with sufficient and appropriate details [51, 11]. These findings motivate our design of PhyT2V that embeds contexts of real-world knowledge and physical rules into T2V prompts, to guide the T2V process for better physical accuracy and temporal alignment. The existing works, however, could still fail when tackling more intricate scenarios such as multi-object interactions, because the T2V model lacks an efficient feedback mechanism to learn how the generated video deviates from the real-world knowledge and physical rules. Researchers suggest to provide such feedback with extra input modalities to T2V models such as sampled video frames, depth map or scribble images [44, 52], but incur significant amounts of extra computing overhead and cannot be generalizable. Instead, in our design of PhyT2V, we aim to fully automate the feedback with only text prompts, and enable iterative feedback for the optimum video quality.
@@ 2.2 Using LLM in T2V Generation
LLMs with strong capabilities in natural language processing (NLP) have been a natural choice for prompt refinement in text-to-image and text-to-video generation, and existing work utilized LLMs to interpret text prompts and orchestrate the initial layout configurations [23, 24, 25, 54, 46, 48, 13, 14]. However, since current LLMs lack inherent understandings of the real-world physical laws, using LLMs with simple instructions usually result in videos that appear visually coherent but lack accurate physical realism, particularly when generating scenes with complex object interactions. Furthermore, these approaches frequently rely on static prompts or simple iterative refinements based on bounding box and segmentation map, which may capture basic visual attributes but fail to adapt to nuanced changes that require continuous physical modeling and adjustment.
@@
The Chain-of-Thought (CoT) method is suitable for single data mode processing because it emphasizes linear decomposition and step-by-step reasoning, and it is especially effective for data processing flows that do not require cross-modal synchronization or interaction. However, multimodal data processing involves the fusion of data from different modalities and complex synchronization requirements, which cannot be completed through simple linear decomposition and requires frequent cross-modal information interaction and parallel processing, which exceeds the linear reasoning ability of the CoT method, resulting in limited performance in multimodal tasks. This is the reason why it is hard to directly apply CoT reasoning in T2V process itself as it requires multimodal alignment between the next and video modality. This also motivates us to adopt video captioning and use the video caption in the reasoning process, so that we can conduct CoT and step-back reasoning only in the text domain.
@@ 3 Method
Compared to the previous prompt enhancing methods, PhyT2V’s key contribution is to analyze the semantic mismatch between currently generated video and the prompt, as well as refinements based on such mismatch. Previous methods can be formulated as $p^{\prime}=f_{enhance}(p,\theta)$, where $p$ and $p^{\prime}$ are the original and enhanced prompts, $f_{\text{enhance}}$ is the prompt enhancer, and $\theta$ represents parameters or rules guiding the enhancement. In contrast, PhyT2V further involves the additional information about the T2V process, i.e., $p^{\prime}=f_{\text{enhance}}(p,f_{mismatch}(C(V(p)),p),f_{phy}(p),\theta)$, where $f_{phy}(p)$ analyzes the physical rules to be followed given $p$, $V(p)$ is the currently generated video given prompt $p$, $C$ is the video captioning model and $f_{mismatch}$ finds the semantic mismatch between $C(V(p))$ and $p$. The key advantages are: (1) Semantic awareness: the refinement process explicitly incorporates the semantic mismatch to enable targeted T2V improvements; (2) Physical-world knowledge integration: physical rules derived from $p$ enable guided enhancement; (3) Guided reasoning: unlike prior methods that rely solely on templates or embeddings, PhyT2V dynamically adapts prompt refinement to the semantic mismatch.
@@ 3.1 Prompting in Parallel Subproblems for Local CoT Reasoning
Besides, to ensure proper CoT reasoning, we follow the existing work [40, 22] and provide in-context examples $[E]$ about tasks. To facilitate LLM’s in-context learning [5, 6], $[E]$ is given in the format of QA pairs. That is, instead of fine-tuning a separate LLM checkpoint for each new task, we prompt the LLM with a few input-output exemplars, to demonstrate the task and condition the task’s input-output format to the LLM, to guide the LLM’s reasoning process.
@@ 4 Experiments
  • CogVideoX [51]: It generates 10-second videos from text prompts, with 16 FPS and 768$\times$1360 resolution. It offers two model variants with 2B and 5B parameters.
  • OpenSora 1.2 [53]: As an alternative to OpenAI’s Sora [3], it contains 1.1B parameters and produces videos with 16 seconds, 720p resolution and different aspect ratios.
  • VideoCrafter [4]: With 1.8B parameters, it can generate both images and videos from text prompts, with 576$\times$1024 resolution and a focus on video dynamics.
@@
In addition, we also evaluated PhyT2V using the widely used VBench metrics and benchmarks [15], which allow comprehensive evaluations of the generated video in multiple aspects, including video quality, video-condition consistency, prompt following and human preference alignment.
Baselines: For fair comparison, we only use the existing T2V prompt enhancers as baselines, and other existing work with extra inputs to T2V models [7, 31, 19, 27, 26] are not applicable. We involve two prompt enhancers: 1) Directly using the existing LLM, particularly ChatGPT4, as the prompt enhancer [28, 47]; 2) Promptist [21], which uses reinforcement learning to automatically refine and enhance prompts in the model-preferred way.
@@ 4.1 Improvement of the Generated Video Quality
Our evaluation results on the VBench metrics are shown in Figure 6, where numbers in Round 1 are the T2V model’s original performance in current VBench leaderboard, and iterative prompt refinements by PhyT2V in Round 2 & 3 noticeably improve the performance in many dimensions. In particular, large improvements are noted in most dimensions of Video-Condition Consistency, showing that PhyT2V improves T2V model’s adherence to prompts and real-world physical rules underlying the prompts.
Figure 6: PhyT2V VBench evaluation results with CogVideoX-5B (left) and OpenSora (right)
@@ 4.2 Different Domains of Physical Rules
4.3 Ablation Study
We conduct an ablation study to evaluate the necessity of both the physical rule reasoning (Step 1) and the mismatch reasoning (Step 2) within our PhyT2V workflow, by removing one of these steps from the refinement process to assess its impact on physical-grounded video generation.
@@ 4.3 Ablation Study
Overall, our study shows that both reasoning steps are integral to PhyT2V, contributing to a more robust and semantically-aligned generation of refined prompts in Step 3. More detailed ablation studies are in Appendix C.
@@ 5 Conclusion
In this paper, we present PhyT2V, a novel data-independent T2V generation framework designed to enhance the generalization capability of existing T2V models to out-of-distribution domains. By incorporating CoT reasoning and step-back prompting, PhyT2V systematically refines T2V prompts to ensure adherence to real-world physical principles without necessitating additional model retraining or reliance on additional conditions. Evaluation results indicate that PhyT2V achieves a 2.3x enhancement in physical realism compared to baseline T2V models and outperforms state-of-the-art T2V prompt enhancers by 35%.
Acknowledgments
We thank the reviewers and the area chair for their insightfulcomments and feedback. This work was supported in part by National Science Foundation (NSF) under grant number IIS-2205360, CCF-2217003, CCF-2215042, and National Institutes of Health (NIH) under grant number R01HL170368.
@@ Appendix A Details of Evaluation Setup
In our evaluations, we use 4 T2V generation models, including CogVideoX-5B [50], CogVideoX-2B [50], OpenSora [53] and VideoCrafter [4]. We choose to use these models as they are all built with transformer based diffusion model, which enhanced the semantic adherence by using the cross attention mechanism, and hold a high score on both the VideoPhy dataset and PhyGenBench dataset leader board. We use Tarsier [39] as the video captioning model for it achieves state of art on multiple video question answering datasets, which ensure the precise and detail of the video captioning step in our approach.
Since PhyT2V improves the quality of generated videos through text prompt refinement, we use the Promptist [21] and GPT-4o [47] as the prompt enhancers, with the same model hyper-parameter setting for the baselines to maintain the consistency between the baseline and our approach. The diffusion model generated video length is setted as 6 second and 8 frames per second with frame resolution 720 $\times$ 480.
To fit the maximum token input length of the T2V model, we limit the word length of the refined prompts to 120, by instructing the ChatGPT4 o1-preview model that is used as the LLM for reasoning. To formatting the output and storage in our approach, the ChatGPT4 o1-preview model are instructed to output in JSON format and output results in each step are saved in a CSV file by row. The ChatGPT4 o1-preview is called by using the API with the input prompt separated into system prompt and user prompt to identify the instruction part and the rest of the original prompt for ChatGPT. The first and second step output is embedded to the third step prompt template by replacing the pre-defined place holder.
@@ B.1 Examples of Generated Videos in Different Categories of Physical Rules
In this subsection, we present the additional comparison examples of generated videos of CogVideo-5b on the VideoPhy and PhyGenBench dataset, Fig 11, 12, 13 show the additional comparison of video generation result on fluid to fluid, solid to fluid, solid to solid cases on VideoPhy dataset, Fig 14, 15, 16 show the additional comparison of video generation result on force, optics and thermal cases on PhyGenBench dataset.
B.2 Refinement process details